OpenStudy (logan13):

The value of y varies directly with x2, and y = 64 when x = 4. What is the value of y when x = 6?

OpenStudy (logan13):

@DARTHVADER2900

OpenStudy (welshfella):

as y varies directly with x^2 the relation is y = kx^2 where k is the constant of variation

OpenStudy (logan13):

uh

OpenStudy (welshfella):

to find the value of k plug in the given values of x and y and solve for k here y = 64 and x = 4

OpenStudy (logan13):

so i multiply 60 and 4

OpenStudy (welshfella):

no replace x and y in y = kx^2 so 64 = k * 4^2 solve for k

OpenStudy (logan13):

362.038671968 You can search for any mathematical expression, using functions such as: sin, cos, sqrt, etc. You can find a complete list of functions here. RadDeg x! Inv sin ln π cos log e tan √ Ans EXP xy ( ) % AC 7 8 9 ÷ 4 5 6 × 1 2 3 − 0 . = +

OpenStudy (welshfella):

when you have found k plug this into the relation then you do the last part of the problem.

OpenStudy (logan13):

362.038671968

OpenStudy (logan13):

@welshfella 362

OpenStudy (darthvader2900):

That's right